OpenStudy (anonymous):

PLZ HELP What polynomial has a graph that passes through the given points? (-2, 2) (-1, -1) (1, 5) (3, 67)

OpenStudy (rsadhvika):

you're given options, or supposed to solve this ?

OpenStudy (anonymous):

those r the options

OpenStudy (anonymous):

A. y = -x3 + 4x2 + 2x - 2 B. y = x3 + 4x2 + 2x - 2 C. y = x4 + 4x3 + 2x2 - 2x D. y = x3 - 4x2 - 2x - 2 nm i lied lol

OpenStudy (rsadhvika):

then its easy, pick one of the given points first lets pick (1, 5) plug x=1 in each of the options, and check if it gives u 5

Directrix (directrix):

@britbrat4601 If you are allowed to use technology on the problem, go to this site: http://www.xuru.org/rt/PR.asp#CopyPaste

OpenStudy (anonymous):

ty @Directrix !!

Directrix (directrix):

Glad to help.